OpenStudy (anonymous):

Solve using the relate-to-a-simpler-peoblem strategy. The population of Jamesburg is 12,000 junction city’s population is 3/25 of jamesburg’s. what is junction city’s population? A. 1440 B. 3992 C. 4008 D. 100,000

OpenStudy (anonymous):

Junction = 3/25 * jamesburg J = 3/25 * 12,000